A purse has P218 in P10, P5, and P1 Coins. There are 46 coins in all and there are 4 more P5 coins than the P10 coins. How many of each kind of coin are there?? hope you can help me...
Answer by scott8148(6628)

let x="P10 coins", so x+4="P5 coins" and 46-(2x+4)="P1 coins"

10(x)+5(x+4)+1(46-(2x+4))=218 __ 10x+5x+20+46-2x-4=218 __ 13x+62=218

13x=156